Antonio is building a bookshelf. He has a piece of wood that measures 90 centimeters in length. The directions require that this piece of wood be cut into two pieces, one piece 20 centimeters longer than the other. He uses the equation below to determine where to make the cut.

x + (x+20) = 90

The shorter piece of wood must then be cut in half. Use the equation above to help you determine the length of the shortest piece of wood.

Answer by ReadingBoosters(3246)

Starting with x + (x+20) = 90
Combine like terms
2x + 20 - 20 = 90 - 20
2x = 70
x = 70/2 = 35
...
Once piece is 35 cm
The other piece is 35+20= 55cm
...
The shorter piece is 35 cm
Cutting it in half would yield 2 pieces of 17.5 cm each.
............
Delighted to help.
-Reading Boosters
Wanting for other what we want for ourselves.
www.MyHomeworkAnswers.com